2(2)^2+3(2)(-4)-4(-4)^2=-80
can you show how this was completed

Answer by checkley77(12844)

2(2)^2+3(2)(-4)-4(-4)^2=-80
2*4+3*-8-4*16=-80
8-24-64=-80
8-88=-80
80=-80